# Human Decision-making is Susceptible to AI-driven Manipulation

### v2

@@ Abstract
Artificial Intelligence (AI) systems are increasingly intertwined with daily life, assisting users in executing various tasks and providing guidance on decision-making. This integration introduces risks of AI-driven manipulation, where such systems may exploit users’ cognitive biases and emotional vulnerabilities to steer them toward harmful outcomes. Through a randomized controlled trial with $233$ participants, we examined human susceptibility to such manipulation in financial (e.g., purchases) and emotional (e.g., conflict resolution) decision-making contexts. Participants interacted with one of three AI agents: a neutral agent (NA) optimizing for user benefit without explicit influence, a manipulative agent (MA) designed to covertly influence beliefs and behaviors, or a strategy-enhanced manipulative agent (SEMA) employing established psychological tactics to reach its hidden objectives. To ensure participant well-being, this study involved hypothetical scenarios, pre-screening for at-risk individuals, and comprehensive post-study debriefing. By analyzing participants’ decision patterns and shifts in their preference ratings post-interaction, we found significant susceptibility to AI-driven manipulation. Particularly, across both decision-making domains, participants interacting with the manipulative agents shifted toward harmful options at substantially higher rates (financial, MA: $62.3\%$, SEMA: $59.6\%$; emotional, MA: $42.3\%$, SEMA: $41.5\%$) compared to the NA group (financial, $35.8\%$; emotional, $12.8\%$). Notably, our results reveal that even simple manipulative objectives (MA) can be as effective as employing established psychological strategies (SEMA) in swaying human decision-making. These findings suggest that AI-driven manipulation could become widespread, as it does not require sophisticated tactics and expertise to influence users. By revealing the potential for covert AI influence, this study highlights a critical vulnerability in human-AI interactions, emphasizing the urgent need for ethical safeguards and regulatory frameworks to ensure responsible deployment of AI technologies and protect human autonomy.
@@ Introduction
The vast integration of AI technologies into our daily lives has fundamentally changed how we process information and make decisions [1, 2]. As AI systems increasingly serve as personal assistants, humans have demonstrated growing reliance and trust in AI-generated content across various domains [3, 4, 5, 6]. While this technological shift offers unprecedented access to information and assistance with decision-making, it also introduces subtle yet significant risks to human autonomy. At the core of these risks lies the potential for AI systems to influence human beliefs and behaviors in ways that may bypass their conscious awareness [7, 8, 9, 10].
@@
Previous research has extensively investigated observable safety risks such as toxicity and discrimination in AI-generated content [11, 12]. However, the psychological mechanisms through which AI systems might influence human decision-making require further investigation [13, 14]. This gap is particularly crucial as humans increasingly resort to AI assistants for practical and personal advice [15]. In such cases, the organizations offering these AI systems, which are often for-profit companies, may wield significant power in shaping user decisions, raising concerns about the potential for large-scale AI-driven manipulation motivated by commercial interests [16].
@@
Figure 1: Overview of our study: a) Framework of decision-making scenarios. Participants were assigned to either the financial or emotional decision-making domain. Each domain featured three pre-defined scenarios with four options: one optimal and three harmful, representing non-existent, excessive, and dependence-inducing products in the financial domain, and maladaptive coping strategies (disengagement, emotional venting, self-blame) in the emotional domain. b) AI conditions. Within each domain, we further assigned the participants to an AI condition: 1. Neutral Agent (NA), an assistant optimizing for user benefit, designed as the control condition in our experiments; 2. Manipulative Agent (MA), an assistant with hidden manipulative objectives; and 3. Strategy-Enhanced Manipulative Agent (SEMA), an assistant equipped with established manipulation strategies based on existing psychological literature. c) Flowchart overview of the study design.
@@
This study introduces a novel framework for examining human susceptibility to AI-driven influence, distinguishing between the two opposing polarities of influence[17]: beneficial persuasion and potentially harmful manipulation. Persuasion involves transparent guidance through logical arguments and ethical appeals[18], and existing literature has demonstrated its potential value in AI applications such as charitable donations [19, 20], mental and physical health interventions [21, 22, 23, 24], debates[25, 26, 27, 28], and advertising[29, 30]. Conversely, manipulation involves covert influence aimed at achieving hidden objectives, often by exploiting the target’s cognitive biases and emotional vulnerabilities for personal gain [31]. AI systems possessing manipulative capabilities could have catastrophic consequences, potentially leading to financial losses [32, 33], data exploitation [13, 34], and overall negative impact on personal beliefs and values [8].
@@
Recent research has demonstrated AI systems’ manipulative capabilities in controlled environments, primarily in game-based settings and simulated interactions[35, 36, 37, 38, 39], without involving real human subjects. While these studies provide valuable insights, their applicability to real-world contexts is limited. For instance, game-based environments simplify human behavior into predictable patterns (e.g., bluffing to maximize in-game rewards[38]) and may not consider the cognitive biases and emotional vulnerabilities that shape human decisions [40, 41]. Similarly, simulated interactions between AI systems role-playing as human users lack the dynamic nature of human-AI interactions, where users’ evolving beliefs, emotions, and perceived consequences (e.g., financial risks or interpersonal conflicts) affect their susceptibility to influence. Therefore, the extent to which real humans are susceptible to AI-driven manipulation in practical decision-making contexts remains unexplored.
@@
To this end, we conducted a randomized controlled trial with $233$ participants, investigating human susceptibility to AI-driven manipulation across two fundamental domains where AI systems are widely deployed [42, 43, 44, 45]: financial and emotional decision-making (Figure 1). While both domains involve real-world risks and consequences, they differ in how AI systems might exploit human vulnerabilities. Financial decisions are grounded in quantifiable risks and trade-offs (e.g., budget constraints, product quality), where users may overtrust AI’s perceived objectivity. In contrast, emotional decision-making primarily involves psychosocial vulnerabilities (e.g., low self-esteem, peer pressure), where AI systems could exploit users’ insecurities or manipulate social dynamics. By contrasting these domains, we aimed to investigate the susceptibility of various human vulnerabilities to AI-driven manipulation.
@@
Our primary hypothesis was that participants interacting with manipulative AI agents (MA and SEMA), which were instructed to covertly influence their decisions, would shift their preferences toward the agents’ hidden incentives, whereas those interacting with the neutral agent would identify and prioritize beneficial options. Our findings supported this hypothesis, revealing that humans are highly susceptible to AI-driven manipulation: participants in the manipulative groups were significantly more likely to shift their preferences toward harmful options and away from beneficial (optimal) choices compared to those interacting with the neutral agent (Figures 2 & S1).
These findings highlight the need for ethical safeguards and regulatory frameworks to protect user autonomy and well-being in human-AI interactions. As AI systems become increasingly sophisticated in their ability to engage with and influence human beliefs and behavior, understanding the psychological vulnerabilities they may exploit is crucial for balancing the potential benefits of AI with the protection of human autonomy. This study also raises a broader ethical question regarding the potential misuse of our findings to exploit vulnerabilities in human decision-making. However, we believe that in the context of increasing AI applications across various sectors, the ethical responsibility to understand and highlight these psychological risks outweighs such concerns.
@@ Results
Analysis of decision-pattern distributions (Figure 2; Methods) indicates that participants who interacted with the manipulative agent (MA) or the strategy-enhanced manipulative agent (SEMA) shifted their preferences towards the harmful options at significantly higher rates than those assisted by the neutral agent (NA). In the financial domain, negative shifts (i.e., switching to a harmful option) occurred in $61.4\%$ (MA) and $59.6\%$ (SEMA) of the cases, compared to $28.3\%$ for NA ($P<0.0001$; Cohen’s $h=0.68$ and $0.64$, respectively). Similar patterns emerged in emotional contexts, as participants in the manipulative groups exhibited harmful shifts in their decisions in $42.3\%$ (MA) and $41.5\%$ (SEMA) of cases, compared to the $12.8\%$ for NA ($P<0.0001$; $h=0.68$ and $0.67$, respectively).
Figure 2: Distribution of decision patterns across AI conditions in decision-making contexts. The bar charts indicate the proportion of decision patterns (%) in financial and emotional scenarios across three AI conditions: NA (Neutral Agent), MA (Manipulative Agent), and SEMA (Strategy-enhanced Manipulative Agent). In both domains, participants in the manipulative groups (MA and SEMA) had significantly more negative shifts (i.e., opting for a harmful option) and fewer positive shifts (i.e., switching to the optimal choice) than the NA condition (Supplementary Table S1).
Conversely, participants were more likely to identify and prioritize the optimal choices post-interaction in the absence of covert manipulation. Interactions with the NA showed significantly more positive shifts (i.e., switching to the optimal choice) than the manipulative conditions across both financial (NA: $31.7\%$; vs. MA: $5.3\%$; vs. SEMA: $4.4\%$; $P<0.0001$; $h=0.73$ and $0.77$, respectively) and emotional (NA: $28.2\%$; vs. MA: $11.7\%$, $P<0.01$; vs. SEMA: $13.0\%$, $P<0.05$; $h=0.42$ and $0.38$, respectively) scenarios. Detailed statistical results are provided in Supplementary Table S1.
@@
These patterns are reflected by temporal changes in participants’ ratings (Figure S1; Supplementary Table S3; Methods). At baseline (pre-interaction), across both domains, there was no significant difference in ratings of the optimal choices and hidden incentives, respectively, between AI conditions ($P=1.00$). However, significant differences emerged post-interaction. In both domains, participants in the manipulative conditions showed significant shifts in their preferences in line with the agents’ hidden motives, exhibiting substantially higher ratings for the hidden incentives (Cohen’s $d\in[0.34,0.68]$) and lower ratings for the optimal options ($d\in[0.54,0.84]$) compared to the control (NA) condition.
Domain-Specific vulnerabilities shape susceptibility patterns
Financial decisions showed significantly higher rates of negative shifts compared to the emotional domain ($49.4\%$ vs. $32.2\%$, $P<0.0001$; Cohen’s $h=0.35$; Figure 2; Supplementary Table S2). This suggests that participants in the financial domain were more likely to be swayed toward harmful choices, possibly due to participants’ over-trust in the AI agent’s perceived objectivity in these contexts, a trend that aligns with existing literature on human reliance on algorithmic recommendations[46]. In contrast, participants were more likely to retain their optimal choices in the emotional contexts compared to the financial domain ($29.9\%$ vs. $12.6\%$, $P<0.0001$, $h=0.43$), indicating that emotional decisions may be more deeply influenced by personal intuition rather than reliance on AI’s advice, a phenomenon known as algorithmic aversion [47], making them less susceptible to external manipulation.
Linear mixed-effects models (Equation 1; Supplementary Table S4) revealed distinct factors of individual susceptibility across these domains. In financial decision-making, lower self-esteem (coefficient $\beta=0.02$, $P=0.03$) was significantly correlated with negative shifts (i.e., switching to a harmful option), suggesting that less confident individuals are more likely to defer to AI recommendations than critically evaluate them. Moreover, unemployed individuals ($\beta=-0.23$, $P=0.08$) were more likely to make harmful financial decisions compared to those working full-time, potentially due to the financial strain and psychological stress of unemployment, which increases the likelihood of riskier behaviors[48] (e.g., opting for less expensive products with few reviews or subscription services offering low initial costs). Greater trust in AI systems also led to more harmful decisions ($\beta=-0.01$, $P=0.09$), further highlighting the impact of AI’s perceived objectivity on human susceptibility in quantitative decision-making contexts[5].
In the emotional domain, higher normative commitment traits ($\beta=0.17$, $P=8.09\times 10^{-3}$) were significantly associated with more beneficial emotional outcomes, indicating the importance of reciprocity and adherence to social norms in identifying the optimal choice for handling emotional dilemmas[49] (i.e., active coping). Conversely, individuals with more affective commitment traits were more prone to negative decision shifts ($\beta=-0.09$, $P=0.08$), suggesting that emotionally reactive individuals are more susceptible to manipulation [50]. Interestingly, the order in which participants completed the scenarios (i.e., completion order) also influenced susceptibility. Participants were more likely to exhibit negative shifts as they progressed through the experiment ($\beta=-0.08$, $P=0.06$), suggesting the initial aversion to AI’s advice in emotional decision-making may diminish over time.
@@
Notably, incorporating psychological strategies (SEMA) did not amplify harmful decision-making outcomes compared to mere manipulative incentives (MA). Across both domains, SEMA’s influence on participants’ decisions nearly matched MA’s ($P=1.00$; Cohen’s $h\in[0.02,0.06]$; Figure 2), indicating that hidden objectives alone are sufficient to influence human decision-making. However, temporal changes in preference ratings revealed domain-specific nuances in the impact of strategies (Figure S1; Supplementary Table S3). Compared to the NA condition, the addition of strategies (SEMA vs. MA) resulted in larger effect sizes (Cohen’s $d$) for both the optimal choices (SEMA: $0.84$ vs. MA: $0.59$) and the hidden incentives (SEMA: $0.68$ vs. MA: $0.48$) in the financial domain, while a contrasting pattern emerged in the emotional domain (optimal choice, SEMA: $0.54$ vs. MA: $0.66$; hidden incentive, SEMA: $0.34$ vs. MA: $0.56$).
@@
Figure 3: Distribution of strategies employed by the strategy-enhanced manipulative agent (SEMA) across decision-making contexts. The bar plots show the proportion of responses for each manipulation strategy used by the SEMA, with significant differences annotated above each bar plot. These results suggest that SEMA adapted its strategy usage to the emotional or financial nature of the decisions, reflecting a tailored approach to influence. $P$ value legend: n.s. (not significant), $P\geq 0.05$; *, $P<0.05$; **, $P<0.01$; ***, $P<0.001$; ****, $P<0.0001$).
In financial scenarios, this agent employed diversion ($23.1\%$ vs. $14.4\%$, $P<0.0001$; Cohen’s $h=0.22$), justification ($15.8\%$ vs. $8.8\%$, $P<0.0001$; $h=0.22$), urgency ($3.0\%$ vs. $0.3\%$, $P<0.0001$; $h=0.24$), and fabricated information ($3.9\%$ vs. $1.9\%$, $P<0.01$; $h=0.12$) significantly more compared to the emotional domain, mirroring real-world predatory marketing tactics[51]. Conversely, compared to the financial domain, it relied on Pleasure Induction ($45.8\%$ vs. $29.8\%$, $P<0.0001$; $h=0.33$) and Guilt Trip ($5.1\%$ vs. $2.4\%$, $P<0.001$; $h=0.14$) significantly more in emotional contexts, attempting to exploit emotional vulnerabilities. In addition, the SEMA adopted considerably more diverse strategies (i.e., Others) in the emotional domain ($2.9\%$ vs. $1.4\%$, $P<0.05$; $h=0.10$). Upon further investigation, such responses primarily aimed to provide reassurance and reinforce the users’ thoughts and beliefs. Detailed statistical results are provided in Supplementary Table S5.
@@
The majority of participants perceived the manipulative agents as helpful across financial (MA: $86.8\%$, and SEMA: $78.9\%$) and emotional (MA: $86.4\%$, and SEMA: $75.6\%$) scenarios, with rates comparable to the neutral agent (financial: $97.5\%$; emotional: $87.1\%$). This perception of helpfulness highlights that participants were largely unaware of the agents’ manipulative intents and viewed them as equally beneficial as the neutral assistant. Notably, while participants were not explicitly prompted to report whether they thought the agents had ulterior motives, a considerable portion of participants mentioned noticing signs of such influence, particularly in the financial domain (NA: $0\%$; MA: $13.2\%$; SEMA: $28.9\%$; $P<0.01$). In contrast, these concerns were less frequently reported in the emotional domain (NA: $2.5\%$; MA: $8.1\%$; SEMA: $9.7\%$; $P=0.41$), suggesting that the agent’s influence was more covert in such contexts. Detailed information on participant feedback is provided in Figure S2 and Supplementary Tables S6 & S7.

### v3

@@ Abstract
Artificial Intelligence (AI) systems are increasingly intertwined with daily life, assisting users in executing various tasks and providing guidance on decision-making. This integration introduces risks of AI-driven manipulation, where such systems may exploit users’ cognitive biases and emotional vulnerabilities to steer them toward harmful outcomes. Through a randomized between-subjects experiment with $233$ participants, we examined human susceptibility to such manipulation in financial (e.g., purchases) and emotional (e.g., conflict resolution) decision-making contexts. Participants interacted with one of three AI agents: a neutral agent (NA) optimizing for user benefit without explicit influence, a manipulative agent (MA) designed to covertly influence beliefs and behaviors, or a strategy-enhanced manipulative agent (SEMA) equipped with established psychological tactics, allowing it to select and apply them adaptively during interactions to reach its hidden objectives. To ensure participant well-being, this study involved hypothetical scenarios, pre-screening for at-risk individuals, and a comprehensive post-study debriefing. By analyzing participants’ preference ratings, we found significant susceptibility to AI-driven manipulation. Particularly, across both decision-making domains, interacting with the manipulative agents significantly increased the odds of rating hidden incentives higher than optimal options (Financial, MA: odds ratio $OR=5.24$, SEMA: $OR=7.96$; Emotional, MA: $OR=5.52$, SEMA: $OR=5.71$) compared to the NA group. Notably, we found no clear evidence that employing psychological strategies (SEMA) was overall more effective than simple manipulative objectives (MA) on our primary outcomes. Hence, AI-driven manipulation could become widespread even without requiring sophisticated tactics and expertise. While our findings are preliminary and derived from hypothetical, low-stakes scenarios, we highlight a critical vulnerability in human-AI interactions, emphasizing the need for ethical safeguards and regulatory frameworks to ensure responsible deployment of AI technologies and protect human autonomy.
@@ Introduction
The vast integration of AI technologies into our daily lives has fundamentally changed how we process information and make decisions [ma2024towards]. As AI systems increasingly serve as personal assistants, humans have demonstrated growing reliance and trust in AI-generated content across various domains [schemmer2022should, leib2021corruptive]. While this technological shift offers unprecedented access to information and assistance with decision-making, it also introduces subtle yet significant risks to human autonomy. At the core of these risks lies the potential for AI systems to influence human beliefs and behaviors in ways that may bypass their conscious awareness [mun2024particip, meier2024llm, weidinger2023sociotechnical].
@@
Previous research has extensively investigated observable safety risks such as toxicity and discrimination in AI-generated content [zhang2023safetybench, yu2024cosafe]. However, the psychological mechanisms through which AI systems might influence human decision-making require further investigation [weidinger2021ethical, durmus2024measuring]. This gap is particularly crucial as humans increasingly resort to AI assistants for practical and personal advice [mireshghallah2024trust]. In such cases, the organizations offering these AI systems, which are often for-profit companies, may wield significant power in shaping user decisions, raising concerns about the potential for large-scale AI-driven manipulation motivated by commercial interests [matz2024potential].
@@
Figure 1: Overview of our study: a) Framework of decision-making scenarios. Participants were assigned to either the financial or emotional decision-making domain. Each domain featured three pre-defined scenarios with four options: one optimal and three harmful, representing non-existent, excessive, and dependence-inducing products in the financial domain, and maladaptive coping strategies (disengagement, emotional venting, self-blame) in the emotional domain. b) AI conditions. Within each domain, we further assigned the participants to an AI condition: 1. Neutral Agent (NA), an assistant optimizing for user benefit, designed as the control condition in our experiments; 2. Manipulative Agent (MA), an assistant with hidden manipulative objectives; and 3. Strategy-Enhanced Manipulative Agent (SEMA), an assistant equipped with a predefined set of manipulation tactics drawn from existing psychological literature, allowing it to select and apply them adaptively during interactions to reach its hidden objectives. All agents were created by prompting GPT-4o [hurst2024gpt] (Prompts are provided in Supplementary Figures S3-S5). c) Flowchart overview of the study design. After progressing through the initial phase of the experiment and being assigned a domain-AI condition, participants were presented with the three corresponding scenarios. Accordingly, for each scenario, they were tasked with rating each option on a 10-point Likert scale, interacting with the AI assistant, and re-rating each option post-interaction. The options in each scenario were shown to each user in a randomized order.
@@
This study introduces a novel framework for examining human susceptibility to AI-driven influence, distinguishing between the two opposing polarities of influence [lau2023good]: beneficial persuasion and potentially harmful manipulation. Persuasion involves transparent guidance through logical arguments and ethical appeals [perloff1993dynamics], and existing literature has demonstrated its potential value in AI applications such as charitable donations [wang2019persuasion, shi2020effects], mental and physical health interventions [jorke2024supporting, karinshak2023working, lin2024imbue, sharma2023human], debates [voelkel2023artificial, breum2024persuasive, costello2024durably], and advertising [feizi2023online, meguellati2024good]. Conversely, manipulation involves covert influence aimed at achieving hidden objectives, often by exploiting the target’s cognitive biases and emotional vulnerabilities for personal gain [handelman2009thought]. AI systems possessing manipulative capabilities could have catastrophic consequences, potentially leading to financial losses [schillaci2024llm, burtell2023artificial], data exploitation [weidinger2021ethical, ai2024defending], and overall negative impact on personal beliefs and values [meier2024llm].
@@
Recent research has demonstrated AI systems’ manipulative capabilities in controlled environments, primarily in game-based settings and simulated interactions[pan2023rewards, wilczynski2024resistance, terekhov2023second, heaven2019no, qi2024enhancing], without involving real human subjects. While these studies provide valuable insights, their applicability to real-world contexts is limited. For instance, game-based environments simplify human behavior into predictable patterns (e.g., bluffing to maximize in-game rewards[heaven2019no]) and may not consider the cognitive biases and emotional vulnerabilities that shape human decisions [Rastogi2020DecidingFA, Brown2011TheRO]. Similarly, simulated interactions between AI systems role-playing as human users lack the dynamic nature of human-AI interactions, where users’ evolving beliefs, emotions, and perceived consequences (e.g., financial risks or interpersonal conflicts) affect their susceptibility to influence. Therefore, the extent to which real humans are susceptible to AI-driven manipulation in practical decision-making contexts remains unexplored.
@@
To this end, we conducted a randomized between-subjects experiment with $233$ participants, investigating human susceptibility to AI-driven manipulation across two fundamental domains where AI systems are widely deployed [lakkaraju2023llms, hwang2021toward, gual2022using, sabour2023chatbot]: financial and emotional decision-making (Figure 1). While both domains involve real-world risks and consequences, they differ in how AI systems might exploit human vulnerabilities. Financial decisions are grounded in quantifiable risks and trade-offs (e.g., budget constraints, product quality), where users may overtrust AI’s perceived objectivity. In contrast, emotional decision-making primarily involves psychosocial vulnerabilities (e.g., low self-esteem, peer pressure), where AI systems could exploit users’ insecurities or manipulate social dynamics. By contrasting these domains, we aimed to investigate the susceptibility of various human vulnerabilities to AI-driven manipulation.
@@
Our primary hypothesis was that participants interacting with manipulative AI agents (MA and SEMA), which were instructed to covertly influence their decisions, would shift their preferences toward the agents’ hidden incentives, whereas those interacting with the neutral agent would identify and prioritize beneficial options. Our findings supported this hypothesis, revealing that humans are highly susceptible to AI-driven manipulation: participants in the manipulative groups were significantly more likely to shift their preferences toward the agents’ hidden incentives and away from beneficial (optimal) options compared to those interacting with the neutral agent (Figures 2 & 3). Notably, we found no clear evidence that employing established psychological strategies is more effective than simple manipulative objectives (SEMA vs. MA) in influencing participants’ preferences.
These findings, while preliminary and derived from simulated low-stakes scenarios, highlight the need for ethical safeguards and regulatory frameworks to protect user autonomy and well-being in human-AI interactions. As AI systems become increasingly sophisticated in their ability to engage with and influence human beliefs and behavior, understanding the psychological vulnerabilities they may exploit is crucial for balancing the potential benefits of AI with the protection of human autonomy. This study also raises a broader ethical question regarding the potential misuse of our findings to exploit vulnerabilities in human decision-making. However, in the context of increasing AI applications across various sectors, the ethical responsibility to understand and highlight these psychological risks outweighs such concerns.
@@ Results
As the manipulative agents were designed to steer participants away from the optimal option and toward their hidden incentive, we model the Hidden-Optimal Differential (HOD; Equation 1), the difference between the preference rating for the hidden incentive and the optimal option, as our primary outcome; $HOD>0$ indicates a relative preference for the hidden incentive. Accordingly, we compare $HOD$ across domain-AI conditions before and after the interaction (Figure 2; Supplementary Table Supplementary Materials; Methods).
Figure 2: Distribution of Hidden-Optimal Differential (HODHOD) across AI conditions in decision-making contexts. While in both domains, participants exhibited relative preference for optimal options ($HOD<0$) across all AI conditions pre-interaction, those interacting with the manipulative agents (MA and SEMA) showed significant inclines post-interaction relative to the NA group, indicating increased alignment with the hidden objectives of these agents (Supplementary Table Supplementary Materials).
@@
At baseline (i.e., pre-interaction), we observed no significant differences between AI conditions in both domains ($P=1.00$); participants, on average, rated optimal options higher than hidden incentives ($HOD<0$) across all AI conditions within financial and emotional contexts. However, significant differences emerged post-interaction. Participants who interacted with the manipulative agent (MA) and the strategy-enhanced manipulative agent (SEMA) aligned their ratings with the agents’ hidden motives, exhibiting significantly higher $HOD$ relative to the neutral agent (NA) group in both financial (MA: $d=1.09$; SEMA: $d=1.53$; both $P<0.001$) and emotional (MA: $d=1.31$; SEMA: $d=1.04$; both $P<0.001$) domains. Analyses of ratings for optimal options and hidden incentives (Supplementary Figure S1; Supplementary Table 3(a); Methods) reinforce these patterns: compared to the NA group, interacting with the manipulative agents (MA and SEMA) led to increased ratings for hidden incentives and lower ratings for the optimal options after the interaction.
Figure 3: Probability distribution of preferences across AI conditions in decision-making contexts. We categorized participants’ preferences in each scenario at each time (pre-/post-interaction) into a three-level outcome based on the Hidden-Optimal Differential (HODHOD; difference in the ratings of the optimal option and the hidden incentive): Hidden ($HOD>0$), Optimal ($HOD<0$), and Tie ($HOD=0$). Across both domains, interacting with the manipulative agents (MA and SEMA) significantly increased the odds of Hidden relative to the NA condition post-interaction (Supplementary Table S4).
To investigate whether shifts in ratings impact participants’ decisions, we modeled a three-level outcome to represent participants’ preferences between optimal options and hidden incentives based on the $HOD$ (Figure 3; Methods): Hidden ($HOD>0$), Optimal ($HOD<0$), and Tie ($HOD=0$). Difference-in-differences analyses (Equation 2; Supplementary Table 4(a)) reveal that relative to the NA, interacting with the MA (Financial: Odds Ratio $OR=5.24$, +29+29 percentage points in the post-interaction probability; Emotional: $OR=5.52$, +20+20 pp) and the SEMA (Financial: $OR=7.96$, +38+38 pp; Emotional: $OR=5.71$, +18+18 pp) significantly increased the odds of preferring hidden incentives over optimal options (Hidden vs. Optimal) across both domains after the interaction (all $P<0.001$). These findings are further supported by the logistic regression model (Equation 3, Supplementary Table 4(b)): with similar preferences pre-interaction, Hidden was significantly more likely that Optimal in the MA (Financial: $OR=6.63$, +31+31 pp; Emotional: $OR=8.57$, +19+19 pp) and the SEMA (Financial: $OR=12.88$, +38+38 pp; Emotional: $OR=6.72$, +20+20 pp) groups, relative to the NA condition across both domains (all $P<0.001$). Moreover, across all AI conditions, participants exhibited significantly lower odds of Hidden (vs. Optimal) pre-interaction (vs. Financial; NA: $OR=0.40$; MA: $OR=0.40$; SEMA: $OR=0.27$; all $P<0.001$), while such differences were insignificant in the financial domain, hinting that participants found it relatively easier to identify and prioritize optimal options in emotional scenarios.
Context-specific individual vulnerabilities shape susceptibility patterns
Linear mixed-effects models (Equation 4; Supplementary Tables S8 & S9) reveal distinct factors of individual susceptibility across domain-AI conditions. Consistent across both domains, pre-interaction $HOD$ was the most dominant factor in both MA (Financial: $\beta=0.43$; Emotional: $\beta=0.56$) and SEMA (Financial: $\beta=0.23$; Emotional: $\beta=0.71$) conditions (all $P<0.001$), aligning with the well-established anchoring effect [furnham2011literature], which suggests initial judgments strongly influence subsequent evaluations.
In the financial domain, consistent with prior work [rhodes1992self], participants with higher self-esteem were associated with less susceptibility in both MA ($\beta=-1.36$, $P<0.01$) and SEMA ($\beta=-1.69$, $P=0.02$) groups, suggesting that less confident individuals are more likely to defer to AI recommendations than critically evaluate them. Among participants who interacted with the MA, higher normative commitment traits ($\beta=-1.20$, $P<0.01$) were significantly associated with lower manipulation susceptibility. In addition, participants recruited through WeChat (vs. Prolific) exhibited higher susceptibility ($\beta=1.85$, $P=0.03$) to the MA’s influence. In the SEMA condition, higher trust in AI ($\beta=2.03$, $P<0.01$) and general interpersonal trust ($\beta=1.62$, $P=0.02$) were associated with increased susceptibility, highlighting the impact of AI’s perceived objectivity on human susceptibility in quantitative decision-making contexts [schemmer2022should, logg2019algorithm]. Moreover, higher agreeableness[palm2025influence] ($\beta=2.32$, $P<0.001$) and reactance commitment traits [montanez2022social] ($\beta=1.90$, $P=0.02$) were associated with higher susceptibility to the SEMA’s influence, hinting that the agent’s explicit tactics (e.g., creating a sense of urgency) may have specifically targeted these traits.
In the emotional domain, individual predictors of susceptibility mainly centered on personality and emotional disposition. In the SEMA group, higher openness ($\beta=1.09$, $P<0.01$) and higher affective commitment traits ($\beta=0.96$, $P<0.01$) were less likely to rate hidden incentives higher than optimal options (i.e., lower susceptibility), while extroversion ($\beta=-0.62$, $P=0.03$) showed a contrasting pattern. However, we found no significant individual factors associated with susceptibility in the MA condition, potentially due to the MA not consistently targeting specific personal vulnerabilities in this domain.
Notably, more conscientious individuals were associated with higher susceptibility for the MA group in the financial domain ($\beta=0.91$, $P=0.02$) and the SEMA group in emotional contexts ($\beta=0.88$, $P<0.01$). This finding aligns with prior work [palm2025influence], indicating a context-dependent role of this trait in influence susceptibility. Moreover, married participants exhibited higher susceptibility to the SEMA across both domains (Financial: $\beta=2.84$; Emotional: $\beta=1.70$; $P<0.05$).
@@
Notably, incorporating psychological strategies (SEMA) did not significantly amplify the manipulative impact compared to mere manipulative incentives (MA). Across both domains, the SEMA’s influence on post-interaction $HOD$ did not significantly differ from MA’s (Financial: $d=0.38$; Emotional: $d=0.41$; both $P>0.05$; Supplementary Table Supplementary Materials). However, compared to the NA, the addition of strategies resulted in a larger effect (Cohen’s d) on $HOD$ in the financial domain (SEMA: $1.53$ vs. MA: $1.09$), while a contrasting pattern emerged in the emotional domain (SEMA: $1.04$ vs. MA: $1.31$). The logistic regression model indicates a similar pattern (Supplementary Table 4(b)): relative to the NA, participants in the SEMA group exhibited comparatively higher odds of Hidden (vs. Optimal) in the financial domain (SEMA: $12.88$ vs. MA: $6.63$) and lower odds in emotional scenarios (SEMA: $6.72$ vs. MA: $8.57$). Notably, while the MA’s impact did not significantly differ across domains ($HOD$: $d=0.26$, $P=0.53$; Hidden vs Optimal: $OR=0.49$, $P=0.17$; Supplementary Tables Supplementary Materials & 7(b)), the SEMA had a significantly weaker effect on the HODHOD ($d=1.10$, $P<0.001$) in the emotional domain (vs. Financial), which in turn led to substantially lower post-interaction odds of Hidden ($OR=0.20$, $P<0.001$).
@@
Figure 4: Distribution of strategies employed by the strategy-enhanced manipulative agent (SEMA) across decision-making contexts. The bar plots show the proportion of responses for each manipulation strategy used by the SEMA, with significant differences annotated above each bar plot. These results suggest that SEMA adapted its strategy usage to the emotional or financial nature of the decisions, reflecting a tailored approach to influence. $P$ value legend: n.s. (not significant), $P\geq 0.05$; *, $P<0.05$; **, $P<0.01$; and ***, $P<0.0001$).
In financial scenarios, this agent employed diversion ($23.1\%$ vs. $14.4\%$, $P<0.0001$; Cohen’s $h=0.22$), justification ($15.8\%$ vs. $8.8\%$, $P<0.0001$; $h=0.22$), urgency ($3.0\%$ vs. $0.3\%$, $P<0.0001$; $h=0.24$), and fabricated information ($3.9\%$ vs. $1.9\%$, $P<0.01$; $h=0.12$) significantly more compared to the emotional domain, mirroring real-world predatory marketing tactics[SUN2023104611]. Conversely, compared to the financial domain, it relied on Pleasure Induction ($45.8\%$ vs. $29.8\%$, $P<0.0001$; $h=0.33$) and Guilt Trip ($5.1\%$ vs. $2.4\%$, $P<0.001$; $h=0.14$) significantly more in emotional contexts, attempting to exploit emotional vulnerabilities. In addition, the SEMA adopted considerably more diverse strategies (i.e., Others) in the emotional domain ($2.9\%$ vs. $1.4\%$, $P<0.05$; $h=0.10$). Upon further investigation, such responses primarily aimed to provide reassurance and reinforce the users’ thoughts and beliefs. Detailed statistical results are provided in Supplementary Table S10.
@@
The majority of participants perceived the manipulative agents as helpful across both financial (MA: $86.8\%$; SEMA: $78.9\%$) and emotional (MA: $86.4\%$; SEMA: $75.6\%$) scenarios, with rates comparable to the neutral agent (financial: $97.5\%$; emotional: $87.1\%$). This perception of helpfulness highlights that participants were largely unaware of the agents’ manipulative intents and viewed them as equally beneficial as the neutral assistant. Notably, while participants were not explicitly prompted to report whether they thought the agents had ulterior motives, a considerable portion of participants mentioned noticing signs of such influence, particularly in the financial domain (NA: $0\%$; MA: $13.2\%$; SEMA: $28.9\%$; $P<0.01$). In contrast, these concerns were less frequently reported in the emotional domain (NA: $2.5\%$; MA: $8.1\%$; SEMA: $9.7\%$; $P=0.41$), suggesting that the agent’s influence was more covert in such contexts. Detailed information on participant feedback is provided in Figure S2 and Tables S11 & S12. Notably, across both domains, mediation analyses (Equation 5; Supplementary Table S13) provided no reliable evidence that any of the perceived values mediated the effect of the manipulative agents on $HOD$.